# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ Appendix A Chains on $1$-dimensional manifolds
In this appendix we prove that, in contrast with the results established in Corollaries 7.6 and 7.9, chains on $1$-dimensional manifolds are classifiable by countable structures and have a generic element. We prove these results for $[0,1]$ and then reduce the case for $S^{1}$ to them.
@@ A.1. Chains on $[0,1]$
$\operatorname{Homeo}([0,1])$ acts on $T$ diagonally, that is, $h\cdot(x,y)=(h(x),h(y))$. The induced action on $\Phi([0,1])\subseteq C(T)$ coincides with the natural action $\operatorname{Homeo}([0,1])\curvearrowright\Phi([0,1])$, discussed in Section 2.
@@
Given a chain $\mathcal{C}\subseteq T$, a closed subset $\mathcal{I}\subseteq\mathcal{C}$ is called a subchain if it is connected (in $T$). A non-trivial subchain $\mathcal{I}$ is
@@
A subchain $\mathcal{I}\subseteq\mathcal{C}$ is maximal horizontal if it is horizontal and it is not properly contained in any horizontal subchain. Analogously, we define maximal vertical subchains and maximal strictly monotone subchains. By Zorn, each horizontal/vertical/strictly monotone subchain, is contained in a maximal subchain of the same type. A subchain is maximal if it is either maximal horizontal, maximal vertical, or maximal strictly monotone.
@@
$\mathcal{E}^{H}$$\coloneqq\{(\mathcal{C},(x,y)):(x,y)\text{ is the maximum of a maximal horizontal subchain in }\mathcal{C}\},$
$\mathcal{E}^{V}$$\coloneqq\{(\mathcal{C},(x,y)):(x,y)\text{ is the maximum of a maximal vertical subchain in }\mathcal{C}\},$
$\mathcal{E}^{S}$$\coloneqq\{(\mathcal{C},(x,y)):(x,y)\text{ is the maximum of a maximal strictly monotone subchain in }\mathcal{C}\}.$
@@
$(x,y)\in\mathcal{C}$ is a closed condition. Then $(x,y)$ is the maximum of a horizontal subchain in $\mathcal{C}$ if and only if $(x+1/n,y)\in\mathcal{C}$ for some $n\in\mathbb{N}$, which is a $\mathbf{\Sigma}^{0}_{2}$ condition. Finally, it is the maximum of a maximal horizontal subchain in $\mathcal{C}$ if and only if moreover $(x-1/n,y)\not\in\mathcal{C}$ for all $n\in\mathbb{N}$, which is a $\mathbf{\Pi}^{0}_{2}$ condition, showing that $\mathcal{E}^{H}$ is Borel. The argument for $\mathcal{E}^{V}$ is exactly the same so it remains to show that $\mathcal{E}^{S}$ is Borel. Indeed, $(x,y)$ is the maximum of a maximal strictly monotone subchain in $\mathcal{C}$ if and only if the following conditions are all satisfied:
@@
  • there exists $n\in\mathbb{N}$ such that for all $m\in\mathbb{N}$ there are no $(x_{1},y_{1})$ and $(x_{2},y_{2})\in\mathcal{C}$ such that for $i\in\{1,2\}$, $x\geq x_{i}\geq x+1/n$, $y\leq y_{i}\leq y-1/n$ and either $x_{1}=x_{2}$ and $y_{1}=y_{2}+1/m$ or $y_{1}=y_{2}$ and $x_{1}=x_{2}+1/m$, which is a $\mathbf{\Sigma}^{0}_{3}$ condition,
@@
We are now able to show that there is generic chain on the interval.
@@
We concentrate on $\Phi^{\infty}([0,1])$; the case for chains with finitely many maximal subchains can be shown to be classifiable by countable structures with an analogous, but easier, argument, which we leave to the reader. Since $\Phi^{\infty}([0,1])$ is Borel by Lemma A.3, there is no harm in treating the two cases separately. Let $\mathcal{E}^{\infty}=(\Phi^{\infty}([0,1])\times T)\cap(\mathcal{E}^{H}\cup\mathcal{E}^{V}\cup\mathcal{E}^{S})$, which is Borel by Lemma A.2. Since, for every $\mathcal{C}\in\Phi^{\infty}([0,1])$, the section $\{(x,y)\in T:(\mathcal{C},(x,y))\}\in\mathcal{E}^{\infty}$ is countably infinite, it follows from the Lusin-Novikov uniformization theorem [KecBook]*Exercise 18.12 that there exists a sequence $e_{n}\colon\Phi^{\infty}([0,1])\to T$ of Borel functions such that for every $\mathcal{C}\in\Phi^{\infty}([0,1])$, $\{e_{n}(\mathcal{C}):n\in\mathbb{N}\}$ is an injective enumeration of $\{(x,y)\in T:(\mathcal{C},(x,y))\in\mathcal{E}^{\infty}\}$. In addition to the quantity and type of maximal subchains occurring in $\mathcal{C}$ we only need to code their relative position in a countable structure in order to determine $\mathcal{C}$ up to homeomorphism. Let =$\mathcal{L}=\{P_{H},P_{V},P_{S},<\}$ be the language containing the three unary predicates $P_{H},P_{V}$ and $P_{S}$ together with a binary relation $<$. Given a chain $\mathcal{C}$, consider the structure $\mathcal{M}(\mathcal{C})$ on $\mathbb{N}$ defined as follows:
@@
It remains to verify that two chains $\mathcal{C},\mathcal{C}^{\prime}\in\Phi^{\infty}([0,1])$ are in the same $\operatorname{Homeo}_{+}([0,1])$-orbit if and only if $\mathcal{M}(\mathcal{C})\approx\mathcal{M}(\mathcal{C}^{\prime})$.
@@
Conversely suppose that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are such that the associated structures are isomorphic and fix an isomorphism $\psi\colon\mathcal{M}(\mathcal{C})\to\mathcal{M}(\mathcal{C}^{\prime})$. Let $x_{0}$ and $x_{0}^{\prime}$ be the roots of $\mathcal{C}$ and $\mathcal{C}^{\prime}$ respectively. We enumerate the maximal subchains of $\mathcal{C}$ as $\{\mathcal{I}_{n}\}_{n\in\mathbb{N}}$, where $\mathcal{I}_{n}$ is such that its maximum is $e_{n}(\mathcal{C})$. Analogously we enumerate the maximal subchains of $\mathcal{C}^{\prime}$ as $\{\mathcal{I}^{\prime}_{n}\}_{n\in\mathbb{N}}$ so that $e_{n}(\mathcal{C}^{\prime})$ is the maximum of $\mathcal{I}^{\prime}_{n}$. For all $n$ we define a pair of order-preserving homeomorphisms $h_{n}^{x}\colon\mathrm{dom}(\mathcal{I}_{n})\to\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})$ and $h_{n}^{y}\colon\mathrm{codom}(\mathcal{I}_{n})\to\mathrm{codom}(\mathcal{I}^{\prime}_{\psi(n)})$. We will then verify that $h^{x}=\bigcup_{n}h_{n}^{x}$ is a densely defined, order-preserving partial homeomorphism $[0,x_{0}]\to[0,x_{0}^{\prime}]$ and that $h^{y}$ defined analogously has the same properties as a function $[x_{0},1]\to[x_{0}^{\prime},1]$ so that $h^{x}\cup h^{y}$ can be extended to the desired homeomorphism of $[0,1]$.
Given $n\in\mathcal{M}(\mathcal{C})$, suppose first that $P^{\mathcal{M}(\mathcal{C})}_{H}(n)$ holds, so that $\mathcal{I}_{n}$ is a maximum horizontal subchain of $\mathcal{C}$, with maximum $e_{n}(\mathcal{C})=(x_{1},y_{1})$ and minimum $(x_{2},y_{1})$. Similarly $\mathcal{I}_{\psi(n)}^{\prime}$ is a maximal horizontal subchain of $\mathcal{C}$ with maximum $e_{\psi(n)}(\mathcal{C}^{\prime})=(x_{1}^{\prime},y_{1}^{\prime})$ and minimum $(x_{2}^{\prime},y_{1}^{\prime})$. Notice that if $y_{1}\in[x_{1},x_{2}]$, then $y_{1}=x_{2}$, since $(x_{2},y_{1})\in T$. But, in this case, $n$ is the minimum element of $<^{\mathcal{M}(\mathcal{C})}$, and thus also $\psi(n)$ is the minimum element of $<^{\mathcal{M}(\mathcal{C}^{\prime})}$, so $y_{1}^{\prime}=x_{2}^{\prime}$. In any case, define $h_{n}^{x}\colon\mathrm{dom}(\mathcal{I}_{n})\to\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})$ to be any order preserving homeomorphism and define $h_{n}^{y}(y_{1})\coloneqq y_{1}^{\prime}$. Note that
@@
The case in which $P_{V}^{\mathcal{M}(\mathcal{C})}(n)$ holds is analogous, so it remains to deal with the case in which $P_{S}^{\mathcal{M}(\mathcal{C})}(n)$ holds. By definition this means that $\mathcal{I}_{n}$ is a maximal strictly monotone subchain in $\mathcal{C}$, with maximum $e_{n}(\mathcal{C})=(x_{1},y_{1})$ and minimum $(x_{2},y_{2})$. Similarly $\mathcal{I}^{\prime}_{\psi(n)}$ is a maximal strictly monotone subchain of $\mathcal{C}^{\prime}$ with maximum $e_{\psi(n)}(\mathcal{C}^{\prime})=(x^{\prime}_{1},y^{\prime}_{1})$ and minimum $(x^{\prime}_{2},y^{\prime}_{2})$. Now $\mathcal{I}_{n}$ is the graph of a continuous, strictly decreasing function $f_{1}\colon\mathrm{dom}(\mathcal{I}_{n})\to\mathrm{codom}(\mathcal{I}_{n})$, and similarly we obtain from $\mathcal{I}^{\prime}_{\psi(n)}$ a continuous strictly decreasing function $f_{2}\colon\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})\to\mathrm{codom}(\mathcal{I}^{\prime}_{\psi(n)})$. If $\mathrm{dom}(\mathcal{I}_{n})\cap\mathrm{codom}(\mathcal{I}_{n})\neq\emptyset$, then they meet only in $x_{2}=y_{2}$, since $(x_{2},y_{2})\in T$. As above, in this case, $n$ is the minimum element of $<^{\mathcal{M}(\mathcal{C})}$, and thus also $\psi(n)$ is the minimum element of $<^{\mathcal{M}(\mathcal{C}^{\prime})}$, so $x_{2}^{\prime}=y_{2}^{\prime}$.
@@
(A.2) $\mathcal{I}^{\prime}_{\psi(n)}=\{(x^{\prime},f_{2}(x^{\prime})):x^{\prime}\in\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})\}=\\ =\{(h_{n}^{x}(x),h_{n}^{y}(f_{1}(x))):x\in\mathrm{dom}(\mathcal{I}_{n})\}=\{(h_{n}^{x}(x),h_{n}^{y}(y)):(x,y)\in\mathcal{I}_{n}\}.$
Finally let $h^{x}=\bigcup_{n\in\mathbb{N}}h_{n}^{x}$ and $h^{y}=\bigcup_{n\in\mathbb{N}}h_{n}^{y}$. Note that if $\mathcal{I}_{n}$ and $\mathcal{I}_{m}$ are two maximal subchains of $\mathcal{C}$ whose domains $[x_{1}^{n},x_{2}^{n}]$ and $[x_{1}^{m},x_{2}^{m}]$ have nonempty intersection, then they meet in exactly one point and we must have either $x_{2}^{n}=x_{1}^{m}$ or $x_{1}^{n}=x^{2}_{m}$. Suppose without loss of generality that the former holds, which implies that $m$ is the $<^{\mathcal{M}(\mathcal{C})}$-immediate successor of $n$. We must have that $h_{n}^{x}(x_{2}^{n})$ is the minimum of the domain of $\mathcal{I}^{\prime}_{\psi(n)}$, while $h_{m}^{x}(x_{1}^{m})$ must be the maximum of the domain of $\mathcal{I}^{\prime}_{\psi(m)}$, but since $\psi$ is an isomorphism, $\psi(m)$ is the $<^{\mathcal{M}(\mathcal{C}^{\prime})}$-immediate successor of $\psi(n)$, showing that
@@
that is, $h^{x}$ is well defined. Since $h^{x}$ is the union of a family of injective functions that cohere on their common domains, $h^{x}$ is injective as well. With a similar argument $h^{y}$ is shown to be well defined and injective too.
@@
$h^{x}$ is a densely defined, order-preserving, partial homeomorphism $[0,x_{0}]\to[0,x_{0}^{\prime}]$ with dense image. Similarly $h^{y}$ is a densely defined, order-preserving, partial homeomorphism $[x_{0},1]\to[x_{0}^{\prime},1]$ with dense image.
Proof of claim..

### v3

@@ Appendix A Chains on $1$-dimensional manifolds
In this appendix we prove that, in contrast with the results established in Corollaries 7.6 and 7.9, chains on $1$-dimensional manifolds are classifiable by countable structures and have a generic element. We prove these results for $[0,1]$ and then reduce the case for $S^{1}$ to them. The structure of the proof is inspired by that of [MR2156712]*Section 6.
@@ A.1. Chains on $[0,1]$
$\operatorname{Homeo}_{+}([0,1])$ acts on $T$ diagonally, that is, $h\cdot(x,y)=(h(x),h(y))$. The induced action on $\Phi([0,1])\subseteq C(T)$ coincides with the natural action $\operatorname{Homeo}_{+}([0,1])\curvearrowright\Phi([0,1])$, discussed in Section 2.
@@
Given a chain $\mathcal{C}\subseteq T$, a closed subset $\mathcal{I}\subseteq\mathcal{C}$ is called a subchain (of $\mathcal{C}$) if it is connected (in $T$). A non-trivial subchain $\mathcal{I}$ is
  1. (1)
@@
A subchain $\mathcal{I}\subseteq\mathcal{C}$ is maximal horizontal if it is horizontal and it is not properly contained in any horizontal subchain of $\mathcal{C}$. Analogously, we define maximal vertical subchains and maximal strictly monotone subchains. By Zorn’s Lemma, each horizontal/vertical/strictly monotone subchain, is contained in a maximal subchain of the same type. A subchain is maximal if it is either maximal horizontal, maximal vertical, or maximal strictly monotone.
@@
$\begin{split}\mathcal{E}^{H}\coloneqq\{(\mathcal{C},(x,y)):&(x,y)\text{ is the maximum of a maximal}\\ &\text{horizontal subchain in }\mathcal{C}\},\end{split}$
$\begin{split}\mathcal{E}^{V}\coloneqq\{(\mathcal{C},(x,y)):&(x,y)\text{ is the maximum of a maximal}\\ &\text{vertical subchain in }\mathcal{C}\},\end{split}$
$\begin{split}\mathcal{E}^{S}\coloneqq\{(\mathcal{C},(x,y)):&(x,y)\text{ is the maximum of a maximal}\\ &\text{strictly monotone subchain in }\mathcal{C}\}.\end{split}$
@@
$(x,y)\in\mathcal{C}$ is a closed condition. Then $(x,y)$ is the maximum of a horizontal subchain in $\mathcal{C}$ if and only if $(x+1/n,y)\in\mathcal{C}$ for some $n\in\mathbb{N}$, which is a $\mathbf{\Sigma}^{0}_{2}$ condition. Finally, it is the maximum of a maximal horizontal subchain in $\mathcal{C}$ if and only if moreover $(x-1/n,y)\not\in\mathcal{C}$ for all $n\in\mathbb{N}$, which is a $\mathbf{\Pi}^{0}_{2}$ condition, showing that $\mathcal{E}^{H}$ is Borel. The argument for $\mathcal{E}^{V}$ is exactly the same, so it remains to show that $\mathcal{E}^{S}$ is Borel. Indeed, $(x,y)$ is the maximum of a maximal strictly monotone subchain in $\mathcal{C}$ if and only if the following conditions are all satisfied:
  1. (a)
@@
  2. (b)
    there exists $n\in\mathbb{N}$ such that for all $m\in\mathbb{N}$ there are no $(x_{1},y_{1})$ and $(x_{2},y_{2})\in\mathcal{C}$ such that for $i\in\{1,2\}$, $x\leq x_{i}\leq x+1/n$, $y\geq y_{i}\geq y-1/n$ and either $x_{1}=x_{2}$ and $y_{1}=y_{2}+1/m$ or $y_{1}=y_{2}$ and $x_{1}=x_{2}+1/m$, which is a $\mathbf{\Sigma}^{0}_{3}$ condition,
  3. (c)
@@
We are now able to show that there is a generic chain on the interval.
@@
We concentrate on $\Phi^{\infty}([0,1])$; the case for chains with finitely many maximal subchains can be shown to be classifiable by countable structures with an analogous, but easier, argument, which we leave to the reader. Since $\Phi^{\infty}([0,1])$ is Borel by Lemma A.3, there is no harm in treating the two cases separately. Let $\mathcal{E}^{\infty}\coloneqq(\Phi^{\infty}([0,1])\times T)\cap(\mathcal{E}^{H}\cup\mathcal{E}^{V}\cup\mathcal{E}^{S})$, which is Borel by Lemma A.2. Since, for every $\mathcal{C}\in\Phi^{\infty}([0,1])$, the section $\{(x,y)\in T:(\mathcal{C},(x,y))\in\mathcal{E}^{\infty}\}$ is countably infinite, it follows from the Lusin-Novikov uniformization theorem [KecBook]*Exercise 18.12 that there exists a sequence $e_{n}\colon\Phi^{\infty}([0,1])\to T$ of Borel functions such that for every $\mathcal{C}\in\Phi^{\infty}([0,1])$, $\{e_{n}(\mathcal{C}):n\in\mathbb{N}\}$ is an injective enumeration of $\{(x,y)\in T:(\mathcal{C},(x,y))\in\mathcal{E}^{\infty}\}$. In addition to the quantity and type of maximal subchains occurring in $\mathcal{C}$ we only need to code their relative position in a countable structure in order to determine $\mathcal{C}$ up to homeomorphism. Let $\mathcal{L}\coloneqq\{P_{H},P_{V},P_{S},<\}$ be the language containing the three unary predicates $P_{H},P_{V}$ and $P_{S}$ together with a binary relation $<$. Given a chain $\mathcal{C}$, consider the structure $\mathcal{M}(\mathcal{C})$ on $\mathbb{N}$ defined as follows:
  1. (1)
@@
It remains to verify that two chains $\mathcal{C},\mathcal{C}^{\prime}\in\Phi^{\infty}([0,1])$ are in the same $\operatorname{Homeo}_{+}([0,1])$-orbit if and only if $\mathcal{M}(\mathcal{C})\cong\mathcal{M}(\mathcal{C}^{\prime})$.
@@
Conversely, suppose that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are such that the associated structures are isomorphic and fix an isomorphism $\psi\colon\mathcal{M}(\mathcal{C})\to\mathcal{M}(\mathcal{C}^{\prime})$. Let $x_{0}$ and $x_{0}^{\prime}$ be the roots of $\mathcal{C}$ and $\mathcal{C}^{\prime}$ respectively. We enumerate the maximal subchains of $\mathcal{C}$ as $\{\mathcal{I}_{n}\}_{n\in\mathbb{N}}$, where $\mathcal{I}_{n}$ is such that its maximum is $e_{n}(\mathcal{C})$. Analogously we enumerate the maximal subchains of $\mathcal{C}^{\prime}$ as $\{\mathcal{I}^{\prime}_{n}\}_{n\in\mathbb{N}}$ so that $e_{n}(\mathcal{C}^{\prime})$ is the maximum of $\mathcal{I}^{\prime}_{n}$. For all $n$ we define a pair of order-preserving homeomorphisms $h_{n}^{x}\colon\mathrm{dom}(\mathcal{I}_{n})\to\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})$ and $h_{n}^{y}\colon\mathrm{codom}(\mathcal{I}_{n})\to\mathrm{codom}(\mathcal{I}^{\prime}_{\psi(n)})$. We will then verify that $h^{x}=\bigcup_{n}h_{n}^{x}$ is a densely defined, order-preserving partial homeomorphism $[0,x_{0}]\to[0,x_{0}^{\prime}]$ and that $h^{y}$ defined analogously has the same properties as a function $[x_{0},1]\to[x_{0}^{\prime},1]$ so that $h^{x}\cup h^{y}$ can be extended to the desired homeomorphism of $[0,1]$.
Given $n\in\mathcal{M}(\mathcal{C})$, suppose first that $P^{\mathcal{M}(\mathcal{C})}_{H}(n)$ holds, so that $\mathcal{I}_{n}$ is a maximum horizontal subchain of $\mathcal{C}$, with maximum $e_{n}(\mathcal{C})=(x_{1},y_{1})$ and minimum $(x_{2},y_{1})$. Similarly, $\mathcal{I}_{\psi(n)}^{\prime}$ is a maximal horizontal subchain of $\mathcal{C}$ with maximum $e_{\psi(n)}(\mathcal{C}^{\prime})=(x_{1}^{\prime},y_{1}^{\prime})$ and minimum $(x_{2}^{\prime},y_{1}^{\prime})$. Notice that if $y_{1}\in[x_{1},x_{2}]$, then $y_{1}=x_{2}$, since $(x_{2},y_{1})\in T$. But, in this case, $n$ is the minimum element of $<^{\mathcal{M}(\mathcal{C})}$, and thus also $\psi(n)$ is the minimum element of $<^{\mathcal{M}(\mathcal{C}^{\prime})}$, so $y_{1}^{\prime}=x_{2}^{\prime}$. In any case, define $h_{n}^{x}\colon\mathrm{dom}(\mathcal{I}_{n})\to\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})$ to be any order preserving homeomorphism and define $h_{n}^{y}(y_{1})\coloneqq y_{1}^{\prime}$. Note that
@@
The case in which $P_{V}^{\mathcal{M}(\mathcal{C})}(n)$ holds is analogous, so it remains to deal with the case in which $P_{S}^{\mathcal{M}(\mathcal{C})}(n)$ holds. By definition this means that $\mathcal{I}_{n}$ is a maximal strictly monotone subchain in $\mathcal{C}$, with maximum $e_{n}(\mathcal{C})=(x_{1},y_{1})$ and minimum $(x_{2},y_{2})$. Similarly, $\mathcal{I}^{\prime}_{\psi(n)}$ is a maximal strictly monotone subchain of $\mathcal{C}^{\prime}$ with maximum $e_{\psi(n)}(\mathcal{C}^{\prime})=(x^{\prime}_{1},y^{\prime}_{1})$ and minimum $(x^{\prime}_{2},y^{\prime}_{2})$. Now $\mathcal{I}_{n}$ is the graph of a continuous, strictly decreasing function $f_{1}\colon\mathrm{dom}(\mathcal{I}_{n})\to\mathrm{codom}(\mathcal{I}_{n})$, and similarly we obtain from $\mathcal{I}^{\prime}_{\psi(n)}$ a continuous strictly decreasing function $f_{2}\colon\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})\to\mathrm{codom}(\mathcal{I}^{\prime}_{\psi(n)})$. If $\mathrm{dom}(\mathcal{I}_{n})\cap\mathrm{codom}(\mathcal{I}_{n})\neq\emptyset$, then they meet only in $x_{2}=y_{2}$, since $(x_{2},y_{2})\in T$. As above, in this case, $n$ is the minimum element of $<^{\mathcal{M}(\mathcal{C})}$, and thus also $\psi(n)$ is the minimum element of $<^{\mathcal{M}(\mathcal{C}^{\prime})}$, so $x_{2}^{\prime}=y_{2}^{\prime}$.
@@
(A.2) $\mathcal{I}^{\prime}_{\psi(n)}$ =$=\{(x^{\prime},f_{2}(x^{\prime})):x^{\prime}\in\mathrm{dom}(\mathcal{I}^{\prime}_{\psi(n)})\}$
=$=\{(h_{n}^{x}(x),h_{n}^{y}(f_{1}(x))):x\in\mathrm{dom}(\mathcal{I}_{n})\}$
=$=\{(h_{n}^{x}(x),h_{n}^{y}(y)):(x,y)\in\mathcal{I}_{n}\}.$
Finally, let $h^{x}=\bigcup_{n\in\mathbb{N}}h_{n}^{x}$ and $h^{y}=\bigcup_{n\in\mathbb{N}}h_{n}^{y}$. Note that if $\mathcal{I}_{n}$ and $\mathcal{I}_{m}$ are two maximal subchains of $\mathcal{C}$ whose domains $[x_{1}^{n},x_{2}^{n}]$ and $[x_{1}^{m},x_{2}^{m}]$ have nonempty intersection, then they meet in exactly one point, and we must have either $x_{2}^{n}=x_{1}^{m}$ or $x_{1}^{n}=x_{2}^{m}$. Suppose without loss of generality that the former holds, which implies that $m$ is the $<^{\mathcal{M}(\mathcal{C})}$-immediate successor of $n$. We must have that $h_{n}^{x}(x_{2}^{n})$ is the minimum of the domain of $\mathcal{I}^{\prime}_{\psi(n)}$, while $h_{m}^{x}(x_{1}^{m})$ must be the maximum of the domain of $\mathcal{I}^{\prime}_{\psi(m)}$, but since $\psi$ is an isomorphism, $\psi(m)$ is the $<^{\mathcal{M}(\mathcal{C}^{\prime})}$-immediate successor of $\psi(n)$, showing that
@@
that is, $h^{x}$ is well-defined. Since $h^{x}$ is the union of a family of injective functions that cohere on their common domains, $h^{x}$ is injective as well. With a similar argument $h^{y}$ is shown to be well-defined and injective too.
@@
$h^{x}$ is a densely defined, order-preserving, partial homeomorphism $[0,x_{0}]\to[0,x_{0}^{\prime}]$ with dense image. Similarly, $h^{y}$ is a densely defined, order-preserving, partial homeomorphism $[x_{0},1]\to[x_{0}^{\prime},1]$ with dense image.
Proof of the Claim..